OpenStudy (anonymous):

Help! how do I solve this? Solve and check. r/6=5 2/3

OpenStudy (anonymous):

Can you write this using the equation button, and are we solving for r?

OpenStudy (anonymous):

um i think we are solving for r...

OpenStudy (anonymous):

these are the answer choices: A. 34 B. 30 2/3 C. 20 D.17/18

OpenStudy (anonymous):

alright, The one question I have is in the second part, is that \[5\frac{ 2 }{ 3 } \] ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Alright, first let's convert it into just a fraction, no whole numbers.\[\frac{ 17 }{ 3 } = \frac{ r }{ 6 }\]

OpenStudy (anonymous):

Now we have a fraction equal to a fraction. Do you know cross multiplication?

OpenStudy (anonymous):

umm i think so

OpenStudy (anonymous):

Show me! I want you to be able to solve this problem in the future, like on a test.

OpenStudy (anonymous):

And if you don't, that's okay, I can teach you.